Hi,
Can you help me with this math problem?
13.) Solve the system.
4x2 - 4y2 = 20
3x2 + 2y2 = 35
The answer should be {(3, 2), (-3, 2), (3, -2), (-3, -2)}

Answer by oscargut(2103)

4x2 - 4y2 = 20
3x2 + 2y2 = 35
then
x^2-y^2=5 then x^2=y^2+5
using 2nd eq
3(y^2+5)+2y^2=35
5y^2=20
y^2=4
then
y=2 or y = -2
if y=2 then x^2=9 then x=3 or x=-3
if y=-2 then x^2=9 then x=3 or x=-3
so solutions are
{(3, 2), (-3, 2), (3, -2), (-3, -2)}